# Conditional regression for the Nonlinear Single-Variable Model

### v1

@@ 1.1 The Nonlinear Single-Variable Model
where the underlying curve $\gamma:[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{d}$ is parametrized by arc-length and has length $\mathrm{len}_{\gamma}$ , and the link function $f\,:\,[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{1}$ depends only on one variable. The random vector $X$ is supported in some domain $\Omega_{\gamma}\subseteq\mathbb{R}^{d}$ containing the image of $\gamma$, such that the closest-point projection $\Pi_{\gamma}:\Omega_{\gamma}\to[0,\mathrm{len}_{\gamma}]$, with
@@
with $t=\Pi_{\gamma}X$, $t$ a random variable on $[0,\mathrm{len}_{\gamma}]$ and $Z_{d-1}$ a centered random vector in $\mathbb{R}^{d-1}$. For each unit vector $v\in\mathbb{S}^{d-1}$, $M_{v}\in\mathbb{O}(d)$ is an orthogonal matrix that maps the $d$-th canonical basis vector $\hat{e}_{d}$ to the vector $v$. Therefore, conditioned on $t=t_{0}\in[0,\mathrm{len}_{\gamma}]$, the random vector $M_{\gamma^{\prime}(t_{0})}\begin{pmatrix}Z_{d-1}\\ 0\in\end{pmatrix}\in\operatorname{span}\{\gamma^{\prime}(t_{0})\}^{\perp}$ is the displacement of $X$ away from $\gamma$.
The regression problem for the Nonlinear Single-Variable Model: Given pairs $(X_{i},Y_{i})_{i=1}^{n}$ with $X_{i}$’s independent copies of $X$ as above, and $Y_{i}=F(X_{i})+\zeta_{i}$, with $F$ as in Definition 1 and observational noise $\zeta_{i}$, we goal is to estimate the regression function $F$ on the support of $X$. The $\zeta_{i}$’s are assumed sub-Gaussian, independent among themselves and from the $X_{i}$’s. Note that the distribution of $X$ is unknown, as are both the link function $f$ and the underlying curve $\gamma$. We illustrate one example in Fig.1.
Figure 1: One example of a Nonlinear Single-Variable Model (1): the underlying curve $\gamma$, plotted in black, is a Meyer helix in $\mathbb{R}^{36}$ (details in Appendix C) with $\sigma_{\gamma}=0.5$ and the link function $f\in\mathcal{C}^{0.7}(\mathbb{R}^{1})$ is strictly monotone, and $\zeta\equiv 0$. We generate $n=5000$ samples $X_{i}$ scattered the curve in a tube of radius 6, colored by $Y_{i}=F(X_{i})=f(\Pi_{\gamma}X_{i})$. Left: Random projection of the data onto $\mathbb{R}^{3}$. Right: Orthogonal projection of data onto the first $3$ principal components. The distribution around this curve does not appear to be linearly embeddable in low dimensions without increasing its complexity, see Appendix C.
@@
We ready for an informal version of the main Theorems 3 and 4:
@@ 2 An estimator for the Nonlinear Single-Variable Model
Step 1: extract geometric features of the underlying curve $\gamma$. Given data $\{(X_{i},Y_{i})\}_{i=1}^{n}$, let $R$ be an interval containing all the $Y_{i}$’s, and $\{{R_{l,h}}\}_{h=1}^{l}$ a partition of $R$, either uniform or based on empirical quantiles (so that all ${R_{l,h}}$ contain the same number of points). We partition the data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ into pairs $\{(\widehat{S}_{l,h},\widehat{R}_{l,h})\}_{h=1}^{l}$ where $\widehat{R}_{l,h}:=\{Y_{i}:Y_{i}\in{R_{l,h}}\}$ and $\widehat{S}_{l,h}:=\{X_{i}:Y_{i}\in{R_{l,h}}\}$. Each empirical slice $\widehat{S}_{l,h}$ is the empirical pre-image of the interval $R_{l,h}$ in the output variable $Y$. We perform the Principal Component Analysis (PCA) of each $\widehat{S}_{l,h}$ to obtain its mean ${\widehat{\mu}_{l,h}}$ and its “significant vector” ${\widehat{v}_{l,h}}$: ${\widehat{\mu}_{l,h}}$ will be approximately on $\gamma$, and ${\widehat{v}_{l,h}}$ will be approximately tangent to $\gamma$ at ${\widehat{\mu}_{l,h}}$, yielding a local first-order approximation to $\gamma$.
Step 2: design a distance function ${\mathrm{dist}}(x,h)$ (and its empirical version ${\widehat{\mathrm{dist}}}(x,h)$), based on the geometric shape of ${S_{l,h}}$ (and $\widehat{S}_{l,h}$ respectively), that measure how far the point $x$ is away from the slice ${S_{l,h}}$ ($\widehat{S}_{l,h}$, respectively). By assigning each $x\in\Omega_{\gamma}$ to the “nearest” slice according to this distance function, we partition the domain $\Omega_{\gamma}$ into several local neighborhoods and we use the significant vector ${\widehat{v}_{l,h}}$ to project the points $X_{i}\in\widehat{S}_{l,h}$ onto a one-dimensional interval $I^{(l,h)}$ in each local neighborhood. Because the significant vector ${\widehat{v}_{l,h}}$ is approximately tangential to the curve $\gamma$, this linear projection approximates the nonlinear projection $\Pi_{\gamma}$ in each local neighborhood.
@@ 2.1 Extracting the geometric features of the underlying curve $\gamma$
For each $t\in[0,\mathrm{len}_{\gamma}]$, let $\hat{n}(t):=\frac{\gamma^{\prime\prime}(t)}{\left\|\gamma^{\prime\prime}(t)\right\|}$ be the unit normal vector to $\gamma$, pointing inwards the circle of curvature. For $x\in\Omega_{\gamma}$ we define the signed projected distance from $x$ to $\gamma$ to be $\tilde{d}(x,\gamma):=\langle x-\gamma(\Pi_{\gamma}x),\hat{n}(\Pi_{\gamma}x)\rangle_{\mathbb{R}^{d}}$, which is zero if $x$ is on the curve $\gamma$, positive if $x$ is inside the circle of curvature, and negative otherwise. A direct calculation shows that the compositional structure $F=f\circ\Pi_{\gamma}$ in Definition 1 implies that for $x\in\Omega_{\gamma}$
@@
Therefore, the gradient vectors of points in each level set $\Pi_{\gamma}^{-1}(t_{0})=\{x:\Pi_{\gamma}x=t_{0}\}$ are all parallel to ${\gamma^{\prime}}(t_{0})$, albeit with magnitude depending on the relative position of $x$ and $\gamma$. As a consequence, each $\Pi_{\gamma}^{-1}(t_{0})$ is contained in a hyperplane. If we could perform singular value decomposition for points on each level set $\Pi_{\gamma}^{-1}(t_{0})$, the singular vector corresponding to the 00 singular value would be parallel to ${\gamma^{\prime}}(t_{0})$.
@@
In the Nonlinear Single-Variable Model, we are still going to use empirical slices, but here we cannot perform an aggregation of estimated vectors because the tangent vectors to $\gamma$ are not constant due to the curvature of $\gamma$. Instead, we can only rely on local information to estimate the nearest-point projection onto $\gamma$, which is now a nonlinear function. Each $\widehat{S}_{l,h}$ still approximates some level set $\Pi_{\gamma}^{-1}(t_{0})$ under the Assumption (LCV) below: roughly speaking, when ${R_{l,h}}$ has a sufficiently small diameter, the $\widehat{S}_{l,h}$ is “thin” along the tangential direction $\gamma^{\prime}(t_{0})$ and “wide” along directions orthogonal to $\gamma^{\prime}(t_{0})$. Consequently, PCA on each empirical slice $\widehat{S}_{l,h}$ locally will yield:
@@ 2.3 Local estimator of the link function $f$ and global estimator of the regression function $F$
After assigning a point $x\in\mathbb{R}^{d}$ to an estimated slice $\widehat{S}_{l,{\widehat{h}_{x}}}$ with index ${\widehat{h}_{x}}$, we use piecewise polynomial estimators to regress the link function $f$ on the corresponding slice. As noted above, we also need to consider empirical slices adjacent to $\widehat{S}_{l,{\widehat{h}_{x}}}$. First, we project the data from $\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}$ orthogonally onto the line with direction ${\widehat{v}_{l,\widehat{h}_{x}}}$. We consider the projected data $\{\langle{\widehat{v}_{l,\widehat{h}_{x}}},X_{i}\rangle:X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}\}$, cover it with an interval $I^{(l,{\widehat{h}_{x}})}$, and further partition $I^{(l,{\widehat{h}_{x}})}$ uniformly into $j$ smaller intervals $\{I^{(l,{\widehat{h}_{x}})}_{j,k}\}_{k=1}^{j}$. We use the sample values $\{Y_{i}:X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}\}$ to construct a local polynomial ${\widehat{f}_{j,k|{\widehat{v}_{l,\widehat{h}_{x}}}}}$ on each interval $I^{(l,{\widehat{h}_{x}})}_{j,k}$ of the partition by solving a least squares fitting problem, obtaining a piecewise polynomial estimator ${\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}$ of the link function. The degree of the local polynomials needed to obtain optimal estimation rates depends on the regularity of the function. A proper partition (or scale) is then chosen to minimize the expected mean squared error (MSE) using classical bias-variance trade-off arguments. Composing this local estimator of $f$ with the “nearest” slice assignment gives us a global estimator $F(x)={\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}(\langle{\widehat{v}_{l,\widehat{h}_{x}}},x\rangle)$.
@@
    Compute interval $R$ of range of samples $\{Y_{i}\}_{i=1}^{n}$. Construct $\{{R_{l,h}}\}_{h=1}^{l}$, the uniform partition of $R$ into $l$ intervals whose preimages are $\{\widehat{S}_{l,h}\}_{h=1}^{l}$ where $\widehat{S}_{l,h}=\{X_{i}:Y_{i}\in{R_{l,h}}\}$;
@@
${\widehat{\mu}_{l,h}}=\frac{1}{{n_{l,h}}}\sum_{i}X_{i}\mathbbm{1}\{X_{i}\in\widehat{S}_{l,h}\}$, the empirical mean
${\widehat{\Sigma}_{l,h}}=\frac{1}{{n_{l,h}}}\sum_{i}(X_{i}-{\widehat{\mu}_{l,h}})(X_{i}-{\widehat{\mu}_{l,h}})^{\intercal}\mathbbm{1}\{X_{i}\in\widehat{S}_{l,h}\}$, the empirical covariance matrix
${\widehat{\lambda}_{l,h,m}}$, the $m$-th eigenvalue of ${\widehat{\Sigma}_{l,h}}$, in descending order
${\widehat{v}}_{l,h,m}$, the $m$-th eigenvector of ${\widehat{\Sigma}_{l,h}}$ corresponding to eigenvalue ${\widehat{\lambda}_{l,h,m}}$
$H_{l,h}=\log\left(\frac{({\widehat{\lambda}_{l,h,2}}\times\cdots\times{\widehat{\lambda}_{l,h,d-1}})^{{2}/{(d-2)}}}{{\widehat{\lambda}_{l,h,1}}{\widehat{\lambda}_{l,h,d}}}\right)$
the empirical significant vector ${\widehat{v}_{l,h}}$, equals ${\widehat{v}}_{l,h,d}$ if $H_{l,h}>0$ and to ${\widehat{v}}_{l,h,1}$ otherwise
@@
    Given $x\in\mathbb{R}^{d}$, for each $h\in\mathcal{H}_{l}$ compute the estimated distance between $x$ and $\widehat{S}_{l,h}$
    ${\widehat{\mathrm{dist}}}(x,h)=\begin{cases}|\langle x-{\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}}\rangle|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}}\left\|x-{\widehat{\mu}_{l,h}}\right\|^{2}\ \text{ if }\widehat{H}_{l,h}>0\\ \left\|x-{\widehat{\mu}_{l,h}}\right\|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}}|\langle x-{\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}}\rangle|^{2}\ \ \text{ if }\widehat{H}_{l,h}<0\end{cases}$
@@
    compute the estimated nearest slice index ${\widehat{h}_{x}}=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{l}}{\widehat{\mathrm{dist}}}(x,h)$
@@
    ${\widehat{f}_{j,k|{\widehat{v}_{l,h}}}}=\operatorname*{{arg\,min}}_{\deg(p)\leqslant m}\sum_{X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}}\left|Y_{i}-p(\langle{\widehat{v}_{l,h}},X_{i}\rangle)\right|^{2}\mathbbm{1}_{{I^{(l,h)}_{j,k}}}(\langle{\widehat{v}_{l,h}},X_{i}\rangle)$
@@
    ${\widehat{F}}(x)={\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}(\langle{\widehat{v}_{l,\widehat{h}_{x}}},x\rangle)$

### v2

@@ 1.1 The Nonlinear Single-Variable Model
where the underlying curve $\gamma:[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{d}$ is twice continuously differentiable, is (without loss of generality) parametrized by arc-length and has length $\mathrm{len}_{\gamma}$ , and the link function $f\,:\,[0,\mathrm{len}_{\gamma}]\to\mathbb{R}^{1}$ depends only on one variable. The random vector $X$ is supported in some domain $\Omega_{\gamma}\subseteq\mathbb{R}^{d}$ containing the image of $\gamma$, such that the closest-point projection $\Pi_{\gamma}:\Omega_{\gamma}\to[0,\mathrm{len}_{\gamma}]$, with
@@
with $t=\Pi_{\gamma}X$, $t$ a random variable on $[0,\mathrm{len}_{\gamma}]$ and $Z_{d-1}$ a centered random vector in $\mathbb{R}^{d-1}$. For each unit vector $v\in\mathbb{S}^{d-1}$, we let $M_{v}\in\mathbb{O}(d)\subseteq\mathrm{GL}(d,\mathbb{R})$ denote an orthogonal matrix that maps the $d$-th canonical basis vector $\hat{e}_{d}$ to the vector $v$. Therefore, conditioned on $t=t_{0}\in[0,\mathrm{len}_{\gamma}]$, the random vector $M_{\gamma^{\prime}(t_{0})}\begin{pmatrix}Z_{d-1}\\ 0\end{pmatrix}\in\operatorname{span}\{\gamma^{\prime}(t_{0})\}^{\perp}$ is the displacement of $X$ away from $\gamma$.
The regression problem for the Nonlinear Single-Variable Model: Given pairs $(X_{i},Y_{i})_{i=1}^{n}$ with $X_{i}$’s independent copies of $X$ as above, and $Y_{i}=F(X_{i})+\zeta_{i}$, with $F$ as in Definition 1 and observational noise $\zeta_{i}$, the goal is to estimate the regression function $F$ on the support of $X$. The $\zeta_{i}$’s are assumed sub-Gaussian, independent among themselves and from the $X_{i}$’s. Note that the distribution of $X$ is unknown, as are both the link function $f$ and the underlying curve $\gamma$. We illustrate one example in Fig.1.
Figure 1: One example of a Nonlinear Single-Variable Model (1): the underlying curve $\gamma$, plotted in black, is a Meyer helix in $\mathbb{R}^{36}$ (details in Appendix C) with $\sigma_{\gamma}=0.5$ and the link function $f\in\mathcal{C}^{0.7}(\mathbb{R}^{1})$ is strictly monotone, and $\zeta\equiv 0$. We generate $n=5000$ samples $X_{i}$ scattered near the curve in a tube of radius 6, colored by $Y_{i}=F(X_{i})=f(\Pi_{\gamma}X_{i})$. Left: Random projection of the data onto $\mathbb{R}^{3}$. Right: Orthogonal projection of data onto the first $3$ principal components. The distribution around this curve does not appear to be linearly embeddable in low dimensions without increasing its complexity, see Appendix C.
@@
We are ready for an informal version of the main Theorems 3 and 4:
@@ 2 An estimator for the Nonlinear Single-Variable Model
Step 1: extract geometric features of the underlying curve $\gamma$. Given data $\{(X_{i},Y_{i})\}_{i=1}^{n}$, let $R$ be an interval containing all the $Y_{i}$’s, and $\{{R_{l,h}}\}_{h=1}^{l}$ a partition of $R$, either uniform or based on empirical quantiles (so that all ${R_{l,h}}$ contain the same number of points). We partition the data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ into pairs $\{(\widehat{S}_{l,h},\widehat{R}_{l,h})\}_{h=1}^{l}$ where $\widehat{R}_{l,h}:=\{Y_{i}:Y_{i}\in{R_{l,h}}\}$ and $\widehat{S}_{l,h}:=\{X_{i}:Y_{i}\in{R_{l,h}}\}$. Each empirical slice $\widehat{S}_{l,h}$ is the empirical pre-image of the interval $R_{l,h}$ in the output variable $Y$. For each $l,h$, let ${S_{l,h}}$ denote the conditional distribution $X|Y\in{R_{l,h}}$. Then the set $\widehat{S}_{l,h}$ of sample points is the empirical version of ${S_{l,h}}$, with moments that should approximate well those of ${S_{l,h}}$ provided the number of samples is large enough. We perform the Principal Component Analysis (PCA) of each $\widehat{S}_{l,h}$ to obtain its mean ${\widehat{\mu}_{l,h}}$ and its “significant vector” ${\widehat{v}_{l,h}}$: ${\widehat{\mu}_{l,h}}$ will be approximately on $\gamma$, and ${\widehat{v}_{l,h}}$ will be approximately tangent to $\gamma$ at ${\widehat{\mu}_{l,h}}$, yielding a local first-order approximation to $\gamma$.
Step 2: design a distance function ${\mathrm{dist}}(x,h)$ (and its empirical version ${\widehat{\mathrm{dist}}}(x,h)$), based on the geometric shape of ${S_{l,h}}$ (and $\widehat{S}_{l,h}$ respectively), that measures how far the point $x$ is away from the slice ${S_{l,h}}$ ($\widehat{S}_{l,h}$, respectively). By assigning each $x\in\Omega_{\gamma}$ to the “nearest” slice according to this distance function, we partition the domain $\Omega_{\gamma}$ into several local neighborhoods and we use the significant vector ${\widehat{v}_{l,h}}$ to project the points $X_{i}\in\widehat{S}_{l,h}$ onto a one-dimensional interval $I^{(l,h)}$ in each local neighborhood. Because the significant vector ${\widehat{v}_{l,h}}$ is approximately tangential to the curve $\gamma$, this linear projection approximates the nonlinear projection $\Pi_{\gamma}$ in each local neighborhood.
@@ 2.1 Extracting the geometric features of the underlying curve $\gamma$
For each $t\in[0,\mathrm{len}_{\gamma}]$, let $\hat{n}(t):=\frac{\gamma^{\prime\prime}(t)}{\left\|\gamma^{\prime\prime}(t)\right\|}$ be the unit normal vector to $\gamma$, pointing inwards the circle of curvature. For $x\in\Omega_{\gamma}$ we define the signed projected distance from $x$ to $\gamma$ to be $\tilde{d}(x,\gamma):=\langle x-\gamma(\Pi_{\gamma}x),\hat{n}(\Pi_{\gamma}x)\rangle_{\mathbb{R}^{d}}$, which is zero if $x$ is on the curve $\gamma$, positive if $x$ is inside the circle of curvature, and negative otherwise. A direct calculation shows that the compositional structure $F=f\circ\Pi_{\gamma}$ in Definition 1 implies that, for curve $\gamma$ parameterized by arc-length and for $x\in\Omega_{\gamma}$
@@
Therefore, the gradient vectors of points in each level set $\Pi_{\gamma}^{-1}(t_{0})=\{x:\Pi_{\gamma}x=t_{0}\}$ are all parallel to the unit-vector ${\gamma^{\prime}}(t_{0})$, albeit with magnitude depending on the relative position of $x$ and $\gamma$. As a consequence, each $\Pi_{\gamma}^{-1}(t_{0})$ is contained in a hyperplane. If we could perform singular value decomposition for points on each level set $\Pi_{\gamma}^{-1}(t_{0})$, the singular vector corresponding to the 00 singular value would be parallel to ${\gamma^{\prime}}(t_{0})$.
@@
In the Nonlinear Single-Variable Model, we are still going to use empirical slices, but here we cannot perform an aggregation of estimated vectors because the tangent vectors to $\gamma$ are not constant due to the curvature of $\gamma$. Instead, we can only rely on local information to estimate the nearest-point projection onto $\gamma$, which is now a nonlinear function. Each $\widehat{S}_{l,h}$ still approximates some level set $\Pi_{\gamma}^{-1}(t_{0})$ under the Assumption (LCV) in Section 3 below: roughly speaking, when ${R_{l,h}}$ has a sufficiently small diameter, the $\widehat{S}_{l,h}$ is “thin” along the tangential direction $\gamma^{\prime}(t_{0})$ and “wide” along directions orthogonal to $\gamma^{\prime}(t_{0})$. Consequently, PCA on each empirical slice $\widehat{S}_{l,h}$ locally will yield:
@@ 2.3 Local estimator of the link function $f$ and global estimator of the regression function $F$
After assigning a point $x\in\mathbb{R}^{d}$ to an estimated slice $\widehat{S}_{l,{\widehat{h}_{x}}}$ with index ${\widehat{h}_{x}}$, we use piecewise polynomial estimators to regress the link function $f$ on the corresponding slice. As noted above, we also need to consider empirical slices adjacent to $\widehat{S}_{l,{\widehat{h}_{x}}}$. First, we project the data from $\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}$ orthogonally onto the line with direction ${\widehat{v}_{l,\widehat{h}_{x}}}$. We consider the projected data $\{\langle{\widehat{v}_{l,\widehat{h}_{x}}},X_{i}\rangle:X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}\}$, cover it with an interval $I^{(l,{\widehat{h}_{x}})}$, and further partition $I^{(l,{\widehat{h}_{x}})}$ uniformly into $j$ smaller intervals $\{I^{(l,{\widehat{h}_{x}})}_{j,k}\}_{k=1}^{j}$. We use the sample values $\{Y_{i}:X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}\}$ to construct a local polynomial ${\widehat{f}_{j,k|{\widehat{v}_{l,\widehat{h}_{x}}}}}$ on each interval $I^{(l,{\widehat{h}_{x}})}_{j,k}$ of the partition by solving a least squares fitting problem, obtaining a piecewise polynomial estimator ${\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}$ of the link function. The degree of the local polynomials needed to obtain optimal (up to log factors) estimation rates depends on the regularity of the function. A proper partition (or scale) is then chosen to minimize the expected mean squared error (MSE) using classical bias-variance trade-off arguments. Composing this local estimator of $f$ with the “nearest” slice assignment gives us a global estimator $F(x)={\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}(\langle{\widehat{v}_{l,\widehat{h}_{x}}},x\rangle)$.
@@
    Compute interval $R$ of range of samples $\{Y_{i}\}_{i=1}^{n}$. Construct $\{{R_{l,h}}\}_{h=1}^{l}$, the uniform partition of $R$ into $l$ intervals whose preimages are $\{\widehat{S}_{l,h}\}_{h=1}^{l}$ where $\widehat{S}_{l,h}=\{X_{i}:Y_{i}\in{R_{l,h}}\}$.
@@
${\widehat{\mu}_{l,h}}=\frac{1}{{n_{l,h}}}\sum_{i}X_{i}\mathbbm{1}\{X_{i}\in\widehat{S}_{l,h}\}$, the empirical mean;
${\widehat{\Sigma}_{l,h}}=\frac{1}{{n_{l,h}}}\sum_{i}(X_{i}-{\widehat{\mu}_{l,h}})(X_{i}-{\widehat{\mu}_{l,h}})^{\intercal}\mathbbm{1}\{X_{i}\in\widehat{S}_{l,h}\}$, the empirical covariance matrix;
${\widehat{\lambda}_{l,h,m}}$, the $m$-th eigenvalue of ${\widehat{\Sigma}_{l,h}}$, in descending order;
${\widehat{v}}_{l,h,m}$, the $m$-th eigenvector of ${\widehat{\Sigma}_{l,h}}$ corresponding to eigenvalue ${\widehat{\lambda}_{l,h,m}}$;
$\widehat{H}_{l,h}=\log\left(\frac{({\widehat{\lambda}_{l,h,2}}\times\cdots\times{\widehat{\lambda}_{l,h,d-1}})^{{2}/{(d-2)}}}{{\widehat{\lambda}_{l,h,1}}{\widehat{\lambda}_{l,h,d}}}\right)$  .
Let the empirical significant vector ${\widehat{v}_{l,h}}$ equal ${\widehat{v}}_{l,h,d}$ if $\widehat{H}_{l,h}>0$ and equal ${\widehat{v}}_{l,h,1}$ otherwise.
@@
    Given $x\in\mathbb{R}^{d}$, for each $h\in\mathcal{H}_{l}$, compute the estimated distance between $x$ and $\widehat{S}_{l,h}$:
    ${\widehat{\mathrm{dist}}}(x,h)=\begin{cases}|\langle x-{\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}}\rangle|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}}\left\|x-{\widehat{\mu}_{l,h}}\right\|^{2}\ \text{ if }\widehat{H}_{l,h}>0\\ \left\|x-{\widehat{\mu}_{l,h}}\right\|^{2}+\frac{{\lambda_{d}(\widehat{\Sigma}_{l,h})}}{{\lambda_{1}(\widehat{\Sigma}_{l,h})}}|\langle x-{\widehat{\mu}_{l,h}},{\widehat{v}_{l,h}}\rangle|^{2}\ \ \text{ if }\widehat{H}_{l,h}<0\end{cases}\,.$
@@
    Compute the estimated nearest slice index ${\widehat{h}_{x}}=\operatorname*{{arg\,min}}_{h\in\mathcal{H}_{l}}{\widehat{\mathrm{dist}}}(x,h)$.
@@
    ${\widehat{f}_{j,k|{\widehat{v}_{l,h}}}}=\operatorname*{{arg\,min}}_{\deg(p)\leqslant m}\sum_{X_{i}\in\bigcup_{|h-{\widehat{h}_{x}}|\leqslant 1}\widehat{S}_{l,h}}\left|Y_{i}-p(\langle{\widehat{v}_{l,h}},X_{i}\rangle)\right|^{2}\mathbbm{1}_{{I^{(l,h)}_{j,k}}}(\langle{\widehat{v}_{l,h}},X_{i}\rangle)\ \,.$
@@
    ${\widehat{F}}(x)={\widehat{f}_{j|{\widehat{v}_{l,\widehat{h}_{x}}}}}(\langle{\widehat{v}_{l,\widehat{h}_{x}}},x\rangle)\ \,.$